Lonnie pitches a baseball of mass 0.20 kg. The ball arrives at home plate with a speed of 40 m/s and is batted straight back to Lonnie with a return speed of 60 m/s. If the bat is in contact with the ball for 0.050 s, what is the impulse experienced by the ball?
1
Expert's answer
2020-11-10T06:59:42-0500

We can find the impulse experienced by the ball from the definition of the impulse:


"J=\Delta p=m(v_f-v_i),""J=0.20\ kg\cdot((-60\ \dfrac{m}{s})-40\ \dfrac{m}{s})=-20\ \dfrac{kgm}{s}."

The sign minus means that the impulse experienced by the ball directed toward the pitcher.

Answer:

"J=20\ \dfrac{kgm}{s}," toward the pitcher.